# Assessing the Adversarial Security of Practical Perceptual Hashing Algorithms

### v1

@@ IV-D Experiment Results of Targeted Blackbox Attacks
IV-E Experiment Results of Hash Inversion Attacks
To demonstrate the service provider adversary’s capability to reconstruct original images from their hashes under some contextual knowledge, such as the group of images to which the original images may belong, we conducted experiments using two datasets MNIST and CelebA as known contexts. For each hashing algorithm and dataset, we trained the inversion network depicted in Fig. 1 for 50 epochs, with a batch size of 64 and an initial learning rate of 0.005. We utilized the MSE loss function with the AdamW optimizer and the cosine annealing learning rate scheduler.
Results are presented in Table V. Samples of synthesized images along with their original images can be observed in Fig. 4 and Fig. 5. It is evident that the hash leaks sufficient information for image reconstruction when certain contextual information is known, such as whether the image belongs to MNIST digits or CelebA photos. Although no image quality metrics were given in [12], Fig. 5(c) shows that our proposed model had a performance comparable to that of [12].
In Table V, smaller $L_{2}$ and LPIPS mean better image reconstruction, while larger SSIM means better image reconstruction. PhotoDNA has the longest hashes at 1152-bits, PDQ has 256-bit hashes and NeuralHash has 96-bit hashes. Intuitively, this implies that PhotoDNA hashes contain the most information about the original images while NeuralHash hashes contain the least. Table V supported this intuition.
The results in the MNIST for PDQ were left blank because PDQ could not work on MNIST. PDQ outputs quality scores of 0 for every image, and each of the hash values calculated turned out to be a 256-dimensional vector of zeros for MNIST images.
@@ IV-E Experiment Results of Hash Inversion Attacks
TABLE V: Quality of Adversarial Images Created in Hash Inversion Attacks (PH: PhotoDNA. PD: PDQ. NH: NeuralHash)
@@
Figure 4: Samples of true images and adversarial images generated by hash inversion attacks based on hashes of (a) PhotoDNA, and (b) NeuralHash. The numbers are ($L_{2}$, SSIM, LPIPS) measures.
@@
Figure 5: Samples of true images and reconstructed images generated by hash inversion attacks from hashes of PhotoDNA, PDQ, and NeuralHash. The numbers are quality metrics {$L_{2}$, SSIM, LPIPS} of reconstructed images.
V Conclusions
This paper presents a thorough examination of the security of three real-world perceptual hashing algorithms (PHAs): PhotoDNA, PDQ, and NeuralHash, against three distinct types of attacks: standard image editing attacks, malicious adversarial blackbox attacks, and hash inversion attacks. To support the security study, a query-efficient blackbox adversarial attack algorithm named Joint Soft-label and Hard-label Attack (JSHA) is designed to expedite blackbox adversarial attacks while requiring fewer queries. Additionally, the paper proposes a new data-efficient hash inversion network capable of reliably reconstructing original images from all hashes, even those of very short length.
The findings of the study indicate that PHAs exhibit robustness against blackbox adversarial attacks when realistic constraints regarding adversarial image distortion and query budget are taken into consideration. However, the vulnerabilities to certain image editing methods and the potential information leakage through the hashes raise significant security concerns. These insights underscore the importance of further research and development efforts to enhance the resilience and security of perceptual hashing algorithms in practical applications.

### v2

@@ A.4.2 Experiment data of untargeted attacks under proposed defense
A.5.1 Experiment data and sample images of hash-inversion attacks without defense
Besides the experiment sample images presented in Section 4.3 and Fig. 2, a complete set of experiment data of image reconstruction quality are in Table 10. In addition, Fig. 5 shows sample images of MNIST dataset reconstruction, Fig. 6 shows sample images of STL-10 dataset reconstruction, whereas Fig. 7 shows sample images of CelebA dataset reconstruction. All these experiment results were obtained without the defense being applied.
Generally, the hash-inversion attacks worked with some success when there was a lot of regularity in the dataset (e.g., MNIST and CelebA). In this case, the hash-inversion algorithm gave an image of such regularity only. The longer hash values leaked more information and allowed better hash-inversion. Nevertheless, even in the best case, it was still impossible to match the reconstructed human faces with the true faces in the CelebA images.
When the dataset was diverse (e.g., STL-10), we were not able to learn any meaningful inversions. This demonstrated that the PHAs were robust against hash-inversion attacks in practical applications. In addition, PHAs with smaller number of output bits were always preferable because natural images were not very standardized, so the smaller number of bits could help to prevent information leakage/hash inversion.
The reason for [12] to draw the false conclusion of successful hash-inversion was because only CelebA images were used over the long PhotoDNA hashes, and the authors overlooked the fact that the faces can not be discriminated based on the reconstructed images.
@@ A.5.1 Experiment data and sample images of hash-inversion attacks without defense
Table 10: Quality of Adversarial Images Created in Hash Inversion Attacks (PH: PhotoDNA. PD: PDQ. NH: NeuralHash)
@@
Figure 5: Samples of true MNIST images and adversarial images generated by hash inversion attacks based on hashes of (a) PhotoDNA, and (b) NeuralHash. The numbers are ($L_{2}$, SSIM, LPIPS) measures.
(a) Using PhotoDNA Hash bits
(b) Using PDQ Hash bits
(c) Using NeuralHash hash bits
Figure 6: Samples of true STL-10 images and adversarial images generated by hash inversion attacks.
@@
Figure 7: Samples of true CelebA images and reconstructed images generated by hash inversion attacks from hashes of PhotoDNA, PDQ, and NeuralHash. The numbers are quality metrics {$L_{2}$, SSIM, LPIPS} of reconstructed images.
A.5.2 Experiment data hash-inversion attacks under the proposed defense
To mitigate the potential hash inversion attacks over regular images, we experimented with the proposed random hash perturbation defense algorithm. Some sample images were shown in Fig. 2(b) and explained briefly in Section 4.3. Extra experiment data are shown in Tables 11, 12 and 13.
Compare these data with those in Table 10, it is clear that as we increased the level of perturbation to the hash bits, the quality of reconstructed images decreased as expected. Therefore, the random perturbation defense does help to defend against hash-inversion attacks.
Table 11: Quality of Inverted Images when Defended with q = 0.1 (PH: PhotoDNA. PD: PDQ. NH: NeuralHash)
Dataset Hash $L_{2}$ SSIM LPIPS
PH $0.14\pm 0.03$ $0.41\pm 0.00$ $0.56\pm 0.06$
STL-10 PD $0.24\pm 0.05$ $0.26\pm 0.05$ $0.65\pm 0.02$
NH $0.23\pm 0.05$ $0.14\pm 0.07$ $0.67\pm 0.03$
Table 12: Quality of Inverted Images when Defended with q = 0.2 (PH: PhotoDNA. PD: PDQ. NH: NeuralHash)
Dataset Hash $L_{2}$ SSIM LPIPS
PH $0.18\pm 0.04$ $0.26\pm 0.08$ $0.68\pm 0.05$
STL-10 PD $0.23\pm 0.05$ $0.24\pm 0.08$ $0.65\pm 0.04$
NH $0.23\pm 0.05$ $0.17\pm 0.08$ $0.77\pm 0.07$
Table 13: Quality of Inverted Images when Defended with q = 0.3 (PH: PhotoDNA. PD: PDQ. NH: NeuralHash)
Dataset Hash $L_{2}$ SSIM LPIPS
PH $0.20\pm 0.06$ $0.21\pm 0.09$ $0.71\pm 0.05$
STL-10 PD $0.23\pm 0.06$ $0.20\pm 0.08$ $0.68\pm 0.04$
NH $0.23\pm 0.05$ $0.17\pm 0.08$ $0.78\pm 0.04$